# Spectrum Projection Score: Aligning Retrieved Summaries with Reader Models in Retrieval-Augmented Generation

### v1

@@ Abstract
Large Language Models (LLMs) have shown improved generation performance through retrieval-augmented generation (RAG) following the retriever-reader paradigm, which supplements model inputs with externally retrieved knowledge. However, prior work often evaluates RAG holistically, assessing the retriever and reader jointly, making it difficult to isolate the true contribution of retrieval, particularly given the prompt sensitivity of LLMs used as readers. We introduce Spectrum Projection Score (SPS), a lightweight, supervision‑free metric that allows the reader to gauge the semantic alignment of a retrieved summary with its hidden representation by comparing the area formed by generated tokens from the summary, and the principal directions of subspace in the reader and to measure the relevance. Building on SPS we present xCompress, an inference‑time controller framework that dynamically samples, ranks, and compresses retrieval summary candidates. Extensive experiments on five QA benchmarks with four open source LLMs show that SPS not only enhances performance across a range of tasks but also provides a principled perspective on the interaction between retrieval and generation.
@@ 1 Introduction
Large-context Retrieval-Augmented Generation (RAG) has demonstrated promising capabilities in addressing open-domain question answering tasks (wang-etal-2024-searching; JMLR:v24:23-0037). In the standard pipeline, a retriever locates and compresses external evidence with a compressor language model, and a reader generates the final answer from the compressed summary (mialon2023augmented). A central challenge is to evaluate whether a given summary will actually help the reader answer the question (pmlr-v202-shi23a), particularly given the reader’s sensitivity to summary variations.
@@
Therefore, a metric that measures the compatibility of the reader with different input summaries is crucial for better compressed summary generation. Existing measurements, such as token-level perplexity and its long-context variants, or on embedding similarity computed with mean pooling (zhang2025entropyregularized; liu2025nover; chen2024inside), primarily assess how typical a token sequence is under a language model. To demonstrate the effects of these measurements, we scatter all tokens from one summary in the reader’s embedding space with t-SNE (Figure 1). For this summary, we highlight token embeddings with different selection methods: tokens nearest to the mean-pooled vector, tokens chosen by perplexity-based scoring with the highest predictive probability, and tokens that contribute to max pooling as per-dimension maxima. Perplexity and mean pooling favour centrally clustered, low-content tokens (e.g., the, of, “,”); by contrast, max pooling surfaces boundary tokens near the convex hull with substantive meaning (e.g., Jeff, Emma, French). This evidence suggests that representing a sentence by a single centroid or by the most probable tokens fails to capture the shape of the information carried by the sequence.
Given the observation above, we argue that these token-level predictive probability-related embeddings not be able to fully represent the semantic meaning of a text segment, which aims to project text into only one single point in the space and ignore the shape of the distribution. Instead, it should be captured by the collective “area” covered by the embeddings of all tokens in the sequence. However, two primary challenges arise when adopting this “area-base” approach: 1) Defining the Area: Formally defining the semantic “area” in high-dimensional embedding space is non-trivial. The embedding space is vast, and computing structures like the convex hull that encapsulates all token embeddings can be computationally expensive. 2) Measuring the Area: Even if a well-defined area is obtained, measuring its shape and size accurately, which is defined in high-dimensional space, is complex due to potential non-convexity and irregular geometry.
@@ 2 Related Work
Recent work on RAG summary focuses on condensing retrieved content into query-relevant representations, primarily through text-to-text summarisation or text-to-embedding conversion (wang2023learning; li-etal-2024-refiner; ke-etal-2024-bridging). Text-to-text approaches generate concise summaries through models trained to distil knowledge from larger language models (yoon-etal-2024-compact; xu2024recomp). Alternatively, text-to-embedding methods, such as xRAG (cheng2024xrag), directly convert retrieved passages into embeddings, concatenating them with query embeddings before processing by the reader. These approaches often yield suboptimal alignment between compressed contexts and the downstream reader’s internal representations due to inherent discrepancies in model-specific embedding spaces. In contrast, our framework xCompress introduces an inference-time, training-free strategy that aligns retrieval summaries with the reader’s semantic space via SPS.
@@
Existing evaluation methods for retrieval-based generation predominantly utilise entropy- or perplexity-based metrics, assessing how well a language model predicts tokens given their preceding context (liu2025nover; yu2025rlpr; xu2025direct). Despite their intuitive appeal, these metrics exhibit fundamental limitations. Primarily, perplexity is sensitive to sequence length (wang2022perplexity), often emphasising predictable but semantically trivial tokens. Consequently, perplexity-based methods inadequately capture the semantic coherence and relevance critical to retrieval-based question-answering (agarwal2024many; li2024long; fang2025what). In contrast, SPS evaluates summaries by aligning their semantic distribution with the reader’s embedding space using max pooling and convex hull theory, thereby emphasising boundary tokens that carry greater semantic relevance over trivial ones.
@@ 3 Preliminary: Summarise retrieval passages to align with the reader
Retrieval-based generation typically follows a retriever–reader pipeline: documents are retrieved, summarised, and then provided to a generative reader for answer production (lewis2020retrieval; izacardunsupervised). While summarisation condenses input and highlights salient content (yoon-etal-2024-compact; chengxrag), prior work has largely overlooked how well a summary aligns with the reader model’s internal representation space. We therefore focus on the following question: how can we measure the quality of retrieved summaries from the perspective of their compatibility with the reader?
@@
However, our analyses (Figure 2a) show a weak association between PPL and downstream QA performance when PPL is used to rank summaries. A likely cause is that PPL is highly sensitive to length and stylistic factors (wang2022perplexity), which becomes particularly problematic in the long-context settings common in retrieval-based generation (agarwal2024many; li2024long). We further evaluate a strengthened likelihood-based metric, LongPPL (fang2025what), which discounts length and emphasises key tokens in extended contexts. Although LongPPL mitigates some issues, it still exhibits limited correlation with QA performance (Figure 2b).
These observations suggest a structural misalignment: token-level log-likelihood primarily captures typicality rather than whether a summary’s salient semantics map onto directions that the reader readily encodes and can exploit for answering the query. Related efforts like SePer (daiseper) seek to assess retrieval utility but rely on human preference signals and do not explicitly model representation-level alignment to task performance.
In contrast to log-likelihood-based metrics like perplexity, we propose the Spectrum Projection Score, a representation-level measure that directly targets summary–reader compatibility. Rather than judging a summary by how “typical” it appears token-by-token, Spectrum Projection Score evaluates how the summary’s salient representation aligns with the reader’s internal geometry: we form a bounder vector from max-pooled token states (to retain boundary, content-bearing features) and assess its alignment to the reader’s principal subspace derived from the model’s parameters or hidden representations. The motivation is that: 1) this bounder vector aims to estimate the Essential Supremum, 2) the bounder vector converges to a distribution-specific property. 3) Thus, this property of consistent estimation allows the boundary vector to be used as a robust tool to determine if the boundary areas of two different generators, for example, the retriever and the reader in this task, are aligned  [^1] Due to the length limitation, we detailed the theoretical analysis, including the formal definition and convergence analysis, in Appendix B.. In general, this combines sentence-level projection with salient token-level cues, yielding a simple, training-free score that better reflects the utility of retrieved summaries for generation (formalised in Section 4.2).
@@ 4.1 Overview and Problem Description
We consider the retriever–reader pipeline (chen-etal-2017-reading; lee-etal-2019-latent) for retrieval-based generation. Given a query $q$ and a corpus $\mathcal{D}$, the retriever returns relevant passages $\mathcal{B}$. The top-$N$ passages are compressed into summaries, either in textual (text-to-text) or embedding (text-to-embedding). Each summary candidate is embedded via the reader’s penultimate-layer representations, which are subsequently max-pooled. A norm-guided filter then determines if further candidate sampling is necessary. If sampling is triggered, additional summary candidates are generated. Each summary is evaluated by the Spectrum Projection Score (SPS), which measures alignment between the summary’s representation and the reader’s principal embedding subspace. Finally, the candidate with the lowest SPS, or the initial summary if sampling is skipped, is provided to the reader LLM for answer generation.
@@ 4.2 Spectrum Projection Score: Measuring Alignment with the Reader LLM
Retrieval-based generation compresses retrieved documents before feeding them into a reader LLM to reduce context length and foreground salient content (lewis2020retrieval; izacardunsupervised; yoon-etal-2024-compact; chengxrag). However, entropy- or perplexity-based evaluations are length-biased and only weakly correlated with downstream performance (wang2022perplexity; agarwal2024many; li2024long; daiseper), as they capture typicality rather than whether a summary’s salient semantics are well represented by the reader’s internal geometry. We therefore seek a representation-level metric that scores a summary by its compatibility with the reader.
@@
Intuitively, this measures how much of the salient summary vector $\mathbf{x}$ is captured by the reader’s principal subspace: $P\mathbf{x}$ is the in-subspace component, and the residual $(I-P)\mathbf{x}$ quantifies what lies outside. Hence, a smaller $\text{SPS}(\mathbf{x})$ indicates stronger alignment between the summary and the reader’s core representational geometry, making the summary easier for the reader to generate.
@@ 4.3 Test-time Sampling with Spectrum Projection Score
Retrieval-based generation typically follows a sequential retriever–reader pipeline (hu2025beyond), where an auxiliary LLM compresses retrieved passages into summaries, either textual (text-to-text) (yoon-etal-2024-compact) or embedding-based (text-to-embedding) (cheng2024xrag), which are then provided to the reader for answer generation. A limitation of this unidirectional flow is that compression is performed without regard to the reader’s internal representational geometry. We instead use the reader’s own representation space to guide compression at test time: we generate a set of summary candidate compressions and score each with Spectrum Projection Score (SPS), selecting the summary that best aligns with the reader.
@@
Embedding-level compression maps retrieved passages (and the query) directly to a summary embedding via a trained projector (cheng2024xrag). Because this mapping is deterministic, it offers no native mechanism for sampling diverse candidates. Inspired by soft reasoning with injected noise (hu2025beyond), we introduce probe-based stochasticity at test time. Concretely, we sample $N$ small Gaussian probe vectors $\{\mathbf{e}_{r}\}_{r=1}^{N}$ and append each to the summary–query embedding before passing the fusion representation through the reader LLM. For each probe, we extract the reader’s penultimate-layer hidden state at the probe position, denoted $\mathbf{h}_{r}$, and compute a simple diversity score following (hu2025beyond):
@@
Sampling multiple candidates improves alignment but is expensive if applied universally. We therefore add a lightweight filter that decides whether further sampling is needed. For the initial summary (text-to-text or text-to-embedding), we compute two proxies in the reader’s latent space: the L2 norm of the mean-pooled representation $L2_{mean}$ (captures overall mass) and the L1 norm of the max-pooled representation $L1_{max}$ (captures salient peaks). Their ratio $L2_{mean}/L1_{max}$ serves as a concentration and stability indicator: higher values indicate that the information follows a more skewed distribution, suggesting that the summary is less likely to benefit from additional sampling. In contrast, lower values reflect a more sparse distribution, where further exploration of the alternatives summary may yield additional value [^2] Detailed theoretical discussion is in Appendix B.. We estimate a threshold on the validation set through full sampling of all data. At inference, if the ratio exceeds the threshold, we accept the initial summary; otherwise, we perform the sampling-and-selection procedure guided by Spectrum Projection Score. This preserves most of the accuracy gains while substantially reducing computation.
@@ 5 Experiment
We evaluate our framework on five retrieval-based QA benchmarks: HotpotQA (yang-etal-2018-hotpotqa), 2WikiMulti-hopQA (2Wiki) (ho-etal-2020-constructing), Natural Questions (NQ) (kwiatkowski-etal-2019-natural), TriviaQA (TQA) (joshi-etal-2017-triviaqa), and Musique. Evaluations are conducted on the development sets, except for TQA, which uses the test set. For NQ, we adopt the original test split with the 21M English Wikipedia dump (karpukhin-etal-2020-dense) as the retrieval corpus. Across all datasets, we follow the data splits and associated document corpora released by kim2024sure and yoon-etal-2024-compact.
@@
Following prior work (chen2024inside; dai2025seper), we use the Area Under the Receiver Operating Characteristic curve (AUROC) and Pearson Correlation Coefficient (PCC) to assess the effectiveness of evaluation metrics. AUROC is widely applied to evaluate the measure of uncertainty estimation (chen2024inside), with higher values indicating better discriminative ability. For retrieval-based generation task performance, we focus on open-domain question answering and report Exact Match (EM) and F1 scores. Following rajpurkar-etal-2016-squad, all predictions and gold answers are normalised by lowercasing and removing punctuation to ensure consistency.
@@
We compare Spectrum Projection Score with perplexity, the most common uncertainty-based evaluation metric for large language model predictions (ren2023outofdistribution), along with its variant, LongPPL, specifically designed to improve performance with long contexts (fang2025what). For retrieval summary compression in retrieval-based generation, we select one recent method from each compression paradigm: the text-to-text method CompAct (yoon-etal-2024-compact) and the text-to-embedding method xRag (cheng2024xrag). Additionally, we evaluate various retrieval strategies:
@@
  • (2) Long-Context LLM Summary, which uses LLMs to summarise retrieved passages before answer generation, following recent practices (yoon-etal-2024-compact).
Backbone models. we utilise four open-source LLMs: LLaMA-3.1-8B-Instruct (grattafiori2024llama), Gemma3-12B-Instruct (team2025gemma), and Qwen3-8B (yang2025qwen3) for the text-to-text paradigm and select Mistral 7B (jiang2024mixtral) for the text-to-embedding paradigm since the reader model in the text-to-embedding method xRag (cheng2024xrag) is specifically trained alongside its retriever projector, we directly adopt its original reader LLM.
@@
For retrieval, we adopt Contriever (izacard2022unsupervised) via the BEIR toolkit (thakur2021beir). Following yoon-etal-2024-compact, we retrieve the top-30 documents for fair comparison. In test-time sampling, we set the temperature to 1.0, apply a repetition penalty of 1.2, and generate five summaries per question to balance diversity and efficiency. For reader LLM generation, we use greedy decoding (temperature = 0.0) to eliminate randomness and ensure reproducibility (sun-etal-2023-chatgpt)  [^3] Prompt details in Appendix A.. For norm-guided filtering, we empirically set the threshold as the top-30% value within the validation set.

### v2

# Beyond Perplexity: Let the Reader Select Retrieval Summaries via Spectrum Projection Score

@@ Abstract
Large Language Models (LLMs) have shown improved generation performance through retrieval-augmented generation (RAG) following the retriever-reader paradigm, which supplements model inputs with externally retrieved knowledge. However, prior work often evaluates RAG holistically, assessing the retriever and reader jointly, making it difficult to isolate the true contribution of retrieval, particularly given the prompt sensitivity of LLMs used as readers. We move beyond perplexity and introduce Spectrum Projection Score (SPS), a lightweight and supervision-free metric that allows the reader to gauge the semantic alignment of a retrieved summary with its hidden representation by comparing the area formed by generated tokens from the summary, and the principal directions of subspace in the reader and to measure the relevance. Building on SPS we present xCompress, an inference‑time controller framework that dynamically samples, ranks, and compresses retrieval summary candidates. Extensive experiments on five QA benchmarks with four open-sourced LLMs show that SPS not only enhances performance across a range of tasks but also provides a principled perspective on the interaction between retrieval and generation.
Codehttps://zhanghao-aaai2026-sps.github.io/AAAI2026-SPS/
@@ 1 Introduction
Large-context Retrieval-Augmented Generation (RAG) has demonstrated promising capabilities in addressing open-domain question answering tasks (Wang et al. 2024; Hu et al. 2024). In the standard pipeline, a retriever locates and compresses external evidence with a compressor language model, and a reader generates the final answer from the compressed summary (Mialon et al. 2023). A central challenge is to evaluate whether a given summary will actually help the reader answer the question (Shi et al. 2023), particularly given the reader’s sensitivity to summary variations.
@@
Therefore, a metric that measures the compatibility of the reader with different input summaries is crucial for better compressed summary generation. Existing measurements, such as token-level perplexity and its long-context variants, or on embedding similarity computed with mean pooling (Zhang et al. 2025; Liu et al. 2025; Chen et al. 2024), primarily assess how typical a token sequence is under a language model. To demonstrate the effects of these measurements, we scatter all tokens from one summary in the reader’s embedding space with t-SNE (Figure 1). For this summary, we highlight token embeddings with different selection methods: tokens nearest to the mean-pooled vector, tokens chosen by perplexity-based scoring with the highest predictive probability, and tokens that contribute to max pooling as per-dimension maxima. Perplexity and mean pooling favour centrally clustered, low-content tokens (e.g., the, of, “,”); by contrast, max pooling surfaces boundary tokens near the convex hull with substantive meaning (e.g., Jeff, Emma, French). This evidence suggests that representing a sentence by a single centroid or by the most probable tokens fails to capture the shape of the information carried by the sequence.
Given the observation above, we argue that these token-level predictive probability-related embeddings not be able to fully represent the semantic meaning of a text segment, which aims to project text into only one single point in the space and ignore the shape of the distribution. Instead, it should be captured by the collective “area” covered by the embeddings of all tokens in the sequence. However, two primary challenges arise when adopting this “area-based” approach: 1) Defining the Area: Formally defining the semantic “area” in high-dimensional embedding space is non-trivial. The embedding space is vast, and computing structures like the convex hull that encapsulates all token embeddings can be computationally expensive. 2) Measuring the Area: Even if a well-defined area is obtained, measuring its shape and size accurately, which is defined in high-dimensional space, is complex due to potential non-convexity and irregular geometry.
@@ 2 Related Work
Recent work on RAG summary focuses on condensing retrieved content into query-relevant representations, primarily through text-to-text summarisation or text-to-embedding conversion (Li et al. 2024; Ke et al. 2024). Text-to-text approaches generate concise summaries through models trained to distil knowledge from larger language models (Yoon et al. 2024). Alternatively, text-to-embedding methods, such as xRAG (Cheng et al. 2024a), directly convert retrieved passages into embeddings, concatenating them with query embeddings before processing by the reader. These approaches often yield suboptimal alignment between compressed contexts and the downstream reader’s internal representations due to inherent discrepancies in model-specific embedding spaces. In contrast, our framework xCompress introduces an inference-time, training-free strategy that aligns retrieval summaries with the reader’s semantic space via SPS.
@@
Existing evaluation methods for retrieval-based generation predominantly utilise entropy- or perplexity-based metrics, assessing how well a language model predicts tokens given their preceding context (Liu et al. 2025; Yu et al. 2025). Despite their intuitive appeal, these metrics exhibit fundamental limitations. Primarily, perplexity is sensitive to sequence length (Wang et al. 2022), often emphasising predictable but semantically trivial tokens. Consequently, perplexity-based methods inadequately capture the semantic coherence and relevance critical to retrieval-based question-answering (Agarwal et al. 2024; Fang et al. 2025). The latest efforts like SePer (Dai et al. 2025a) seek to assess retrieval utility but rely on human preference signals and do not explicitly model representation-level alignment to task performance. In contrast, SPS evaluates summaries by aligning their semantic distribution with the reader’s embedding space using max pooling and convex hull theory, thereby emphasising boundary tokens that carry greater semantic relevance over trivial ones.
@@ 3 Preliminary: Summarise retrieval passages to align with the reader
Retrieval-based generation typically follows a retriever–reader pipeline: documents are retrieved, summarised, and then provided to a generative reader for answer production (Lewis et al. 2020; Izacard et al. 2022a; Qiu et al. 2024). While summarisation condenses input and highlights salient content (Yoon et al. 2024; Cheng et al. 2024b), prior work has largely overlooked how well a summary aligns with the reader model’s internal representation space. We therefore focus on the following question: how can we measure the quality of retrieved summaries from the perspective of their compatibility with the reader?
@@
However, our analyses (Figure 2a) show a weak association between PPL and downstream QA performance when PPL is used to rank summaries(Wang et al. 2022; Agarwal et al. 2024), as well as the alternative implementations such as long PPL (Fang et al. 2025). These observations suggest a structural misalignment: token-level log-likelihood primarily captures typicality rather than whether a summary’s salient semantics map onto directions that the reader readily encodes and can exploit for answering the query.
In contrast, we quantify a summary’s semantics by its coverage in representation space. Ideally, this is the volume of the convex hull of its token embeddings, but computing high-dimensional hulls is prohibitive. We therefore use a bounder vector, which is obtained by elementwise max pooling over token states and forms the minimal axis-aligned enclosure that upper-bounds the token-based convex hull. This summary-level enclosure expands only along informative coordinates as more tokens are added, consistent with the intuition that longer inputs yield greater information. In contrast, low-content tokens (Figure 1) typically lie near the centre of the distribution and therefore leave the boundary unchanged. Thus, the bounder provides a simple, stable proxy for a summary’s salient semantics (formal properties in Appendix B). In summary, in this work, the motivation is that: 1) this bounder vector aims to estimate the Essential Supremum, 2) the bounder vector converges to a distribution specific property. 3) Thus, this property of consistent estimation allows the boundary vector to be used as a robust tool to determine if the boundary areas of two different generators, for example, the retriever and the reader in this task, are aligned. In general, this combines sentence-level projection with salient token-level cues, yielding a simple, training-free score that better reflects the utility of retrieved summaries for generation (formalised in Section 4.2). Building on this, the next section defines Spectrum Projection Score, a representation-level score that measures summary–reader compatibility by projecting the max-pooling based “bounder vector” onto the reader’s principal subspace.
@@ 4.1 Overview and Problem Description
We consider the retriever–reader pipeline (Lee, Chang, and Toutanova 2019) for retrieval-based generation. Given a query $q$ and a corpus $\mathcal{D}$, the retriever returns relevant passages $\mathcal{B}$. The top-$N$ passages are compressed into summaries, either in textual (text-to-text) or embedding (text-to-embedding). Each summary candidate is embedded via the reader’s penultimate-layer representations, which are subsequently max-pooled. A norm-guided filter then determines if further candidate sampling is necessary. If sampling is triggered, additional summary candidates are generated. Each summary is evaluated by the Spectrum Projection Score (SPS), which measures alignment between the summary’s representation and the reader’s principal embedding subspace. Finally, the candidate with the lowest SPS, or the initial summary if sampling is skipped, is provided to the reader LLM for answer generation.
@@ 4.2 Spectrum Projection Score: Measuring Alignment with the Reader LLM
Retrieval-based generation compresses retrieved documents before feeding them into a reader LLM to reduce context length and foreground salient content (Lewis et al. 2020; Izacard et al. 2022a; Yoon et al. 2024; Cheng et al. 2024b). However, entropy- or perplexity-based evaluations are length-biased and only weakly correlated with downstream performance (Wang et al. 2022; Agarwal et al. 2024; Dai et al. 2025a), as they capture typicality rather than whether a summary’s salient semantics are well represented by the reader’s internal geometry. We therefore seek a representation-level metric that scores a summary by its compatibility with the reader.
@@
Intuitively, this measures how much the main component of bounder vector $\mathbf{x}$ is captured by the reader’s principal subspace: $P\mathbf{x}$ is the in-subspace component, and the residual $(I-P)\mathbf{x}$ quantifies what lies outside. Hence, a smaller $\text{SPS}(\mathbf{x})$ indicates stronger alignment between the summary and the reader’s core representational geometry, making the summary easier for the reader to generate.
@@ 4.3 Test-time Sampling with Spectrum Projection Score
Retrieval-based generation typically follows a sequential retriever–reader pipeline (Hu et al. 2025), where an auxiliary LLM compresses retrieved passages into summaries, either textual (text-to-text) (Yoon et al. 2024) or embedding-based (text-to-embedding) (Cheng et al. 2024a), which are then provided to the reader for answer generation. A limitation of this unidirectional flow is that compression is performed without regard to the reader’s internal representational geometry. We instead use the reader’s own representation space to guide compression at test time: we generate a set of summary candidate compressions and score each with Spectrum Projection Score (SPS), selecting the summary that best aligns with the reader.
@@
Embedding-level compression maps retrieved passages (and the query) directly to a summary embedding via a trained projector (Cheng et al. 2024a). Because this mapping is deterministic, it offers no native mechanism for sampling diverse candidates. Inspired by soft reasoning with injected noise (Hu et al. 2025; Zhu et al. 2025), we introduce probe-based stochasticity at test time. Concretely, we sample $N$ small Gaussian probe vectors $\{\mathbf{e}_{r}\}_{r=1}^{N}$ and append each to the summary–query embedding before passing the fusion representation through the reader LLM. For each probe, we extract the reader’s penultimate-layer hidden state at the probe position, denoted $\mathbf{h}_{r}$, and compute a simple diversity score following Hu et al. (2025):
@@
Sampling multiple candidates improves alignment but is expensive if applied universally. We therefore add a lightweight filter that decides whether further sampling is needed. For the initial summary (text-to-text or text-to-embedding), we compute two proxies in the reader’s latent space: the L2 norm of the mean-pooled representation $L2_{mean}$ (captures overall mass) and the L1 norm of the max-pooled representation $L1_{max}$ (captures salient peaks). Their ratio $L2_{mean}/L1_{max}$ serves as a concentration and stability indicator: higher values indicate that the information follows a more skewed distribution, suggesting that the summary is less likely to benefit from additional sampling. In contrast, lower values reflect a more sparse distribution, where further exploration of the alternatives summary may yield additional value (Detailed theoretical discussion is in Appendix B). We estimate the threshold by sampling a small subset of the dataset. In inference, if the ratio exceeds the threshold, we accept the initial summary; otherwise, we perform the sampling-and-selection procedure guided by Spectrum Projection Score. This preserves most of the accuracy gains while substantially reducing computation.
@@ 5 Experiment
We evaluate our framework on five retrieval-based QA benchmarks: HotpotQA (Yang et al. 2018), 2WikiMulti-hopQA (2Wiki) (Ho et al. 2020), Natural Questions (NQ) (Kwiatkowski et al. 2019), TriviaQA (TQA) (Joshi et al. 2017), and Musique (Trivedi et al. 2022). Evaluations are conducted on the development sets, except for TQA, which uses the test set. For NQ, we adopt the original test split with the 21M English Wikipedia dump (Karpukhin et al. 2020) as the retrieval corpus. Across all datasets, we follow the data splits and associated document corpora released by Kim et al. (2024) and Yoon et al. (2024).
@@
Following prior work (Chen et al. 2024; Dai et al. 2025b), we use the Area Under the Receiver Operating Characteristic curve (AUROC) and Pearson Correlation Coefficient (PCC) to assess the effectiveness of evaluation metrics. AUROC is widely applied to evaluate the measure of uncertainty estimation (Chen et al. 2024), with higher values indicating better discriminative ability. For retrieval-based generation task performance, we focus on open-domain question answering and report Exact Match (EM) and F1 scores. Following Rajpurkar et al. (2016), all predictions and gold answers are normalised by lowercasing and removing punctuation to ensure consistency.
@@
We compare Spectrum Projection Score with perplexity, the most common uncertainty-based evaluation metric for large language model predictions (Ren et al. 2023), along with its variant, LongPPL, specifically designed to improve performance with long contexts (Fang et al. 2025). For retrieval summary compression in retrieval-based generation, we select one recent method from each compression paradigm: the text-to-text method CompAct (Yoon et al. 2024) and the text-to-embedding method xRag (Cheng et al. 2024a). Additionally, we evaluate various retrieval strategies:
@@
  • (2) Long-Context LLM Summary, which uses LLMs to summarise retrieved passages before answer generation, following recent practices (Yoon et al. 2024).
Backbone models. we utilise four open-source LLMs: LLaMA-3.1-8B-Instruct (Grattafiori et al. 2024), Gemma3-12B-Instruct (Team et al. 2025), and Qwen3-8B (Yang et al. 2025) for the text-to-text paradigm and select Mistral 7B (Jiang et al. 2024) for the text-to-embedding paradigm since the reader model in the text-to-embedding method xRag (Cheng et al. 2024a) is specifically trained alongside its retriever projector, we directly adopt its original reader LLM.
@@
For retrieval, we adopt Contriever (Izacard et al. 2022b) via the BEIR toolkit (Thakur et al. 2021). Following Yoon et al. (2024), we retrieve the top-30 documents for fair comparison. In test-time sampling, we set the temperature to 1.0, apply a repetition penalty of 1.2, and generate five summaries per question to balance diversity and efficiency. For reader LLM generation, we use greedy decoding (temperature = 0.0) to eliminate randomness and ensure reproducibility (Sun et al. 2023)(Prompt details are in Appendix A). For norm-guided filtering, we empirically set the threshold as the top-30% value within the validation set.